# Extremal Bounds on the Properties of Sigma and Albertson Indices for Non-Decreasing Degree Sequences

### v2

@@ 3 Impact Bounds Among Sigma Index
Let $T$ be tree of order $n$, $\Delta_{\mathscr{A}}$ be the maximum degree of $\mathscr{A}$ and $\Delta_{\mathscr{R}}$ be the maximum degree of $\mathscr{R}$. Then, the lower bound of Sigma index satisfy
$\sigma(T)\geqslant\operatorname{irr}(T)+\left\lfloor\frac{n-2}{a_{r}-t_{m}}\right\rfloor+\Delta(\Delta_{\mathscr{A}}-\Delta_{\mathscr{R}})^{2}.$ (4)
@@
Let $\operatorname{irr}(T)$ be the Albertson index of $T$, $\Delta_{\mathscr{A}}$ be the maximum degree of $\mathscr{A}$ and $\Delta_{\mathscr{R}}$ be the maximum degree of $\mathscr{R}$. Then, clearly $\Delta_{\mathscr{A}}>\Delta_{\mathscr{R}}$ and $\Delta>\Delta_{\mathscr{A}}$. Thus, $(\Delta_{\mathscr{A}}-\Delta_{\mathscr{R}})^{2}<\operatorname{irr}(T)<\Delta(\Delta_{\mathscr{A}}-\Delta_{\mathscr{R}})^{2}$. Then, let each leaf $u_{i}$ be adjacent to a vertex $w_{i}$ with degree $\deg_{T}(w_{i})=d_{w_{i}}$. The total contribution from these edges is:
$\sum_{u_{i}\in N(w_{i})}(d_{w_{i}}-1)(d_{w_{i}}-2).$ (5)
Thus, by considering (5) we have
$\sigma(T)-\operatorname{irr}(T)$ $=\sum_{uv\in E(T)}(\deg_{T}(u)-\deg_{T}(v))^{2}-\sum_{uv\in E(T)}|\deg_{T}(u)-\deg_{T}(v)|$
$=\sum_{uv\in E(T)}\left[(\deg_{T}(u)-\deg_{T}(v))^{2}-|\deg_{T}(u)-\deg_{T}(v)|\right]$
$=\sum_{uv\in E(T)}\left[\deg_{T}(u)^{2}+\deg_{T}(v)^{2}-2\deg_{T}(u)\deg_{T}(v)-|\deg_{T}(u)-\deg_{T}(v)|\right]$
$\geqslant\sum_{u\in N(w_{i})}(d_{w_{i}}-1)(d_{w_{i}}-2)+\sum_{v\in N(w_{i})}(d_{w_{i}}-1)(d_{w_{i}}-2).$
Hence, we have
$\sigma(T)-\operatorname{irr}(T)\geqslant\sum_{uv\in E(T)}\left((\deg_{T}(u)-1)(\deg_{T}(u)-2)+(\deg_{T}(v)-1)(\deg_{T}(v)-2)\right).$ (6)
On the other hand, for the second term, since we observe that $n-2>a_{r}-t_{m}$, and to achieve this difference of $a_{r}-t_{m}<\Delta$, we obtain
$\sigma(T)\geqslant\operatorname{irr}(T)+\left\lfloor\frac{n-2}{a_{r}-t_{m}}\right\rfloor.$ (7)
From relations (6) and (7), considering the quantity $\Delta(\Delta_{\mathscr{A}}-\Delta_{\mathscr{R}})^{2}$ as a constant term, where $(\Delta_{\mathscr{A}}-\Delta_{\mathscr{R}})^{2}<\operatorname{irr}(T)<\Delta(\Delta_{\mathscr{A}}-\Delta_{\mathscr{R}})^{2}$.Thus, the relation (4) remains valid by adding this quantity to relation (7). ∎
The upper bound of Sigma index holds among Proposition 3.4 by considering the maximum degree of $\mathscr{A}$ and the maximum degree of $\mathscr{R}$.
@@
Let $T$ be tree of order $n$, $\Delta_{\mathscr{A}}$ be the maximum degree of $\mathscr{A}$ and $\Delta_{\mathscr{R}}$ be the maximum degree of $\mathscr{R}$. Then, the upper bound of Sigma index satisfy
$\sigma(T)\leqslant\sum_{v\in V(T)}\deg_{T}(v)^{3}+\operatorname{irr}(T)+\left\lfloor\frac{n-2}{a_{r}-t_{m}}\right\rfloor+\Delta(\Delta_{\mathscr{A}}-\Delta_{\mathscr{R}})^{2}.$ (8)
@@
According to Proposition 3.3 we find that $(\Delta_{\mathscr{A}}-\Delta_{\mathscr{R}})^{2}<\operatorname{irr}(T)<\Delta(\Delta_{\mathscr{A}}-\Delta_{\mathscr{R}})^{2}$. Thus,
$\sigma(T)\leqslant\sum_{v\in V(T)}\deg_{T}(v)^{3}+\Delta(\Delta_{\mathscr{A}}-\Delta_{\mathscr{R}})^{2}.$ (9)
Then, by considering $n<\Delta(\Delta_{\mathscr{A}}-\Delta_{\mathscr{R}})^{2}$ we have $\sigma(T)\leqslant(n-\Delta_{\mathscr{A}})^{2}+(n-\Delta_{\mathscr{R}})^{2}$. Thus, the Albertson index satisfy $\operatorname{irr}(T)\geqslant 2(\Delta_{\mathscr{A}}^{2}-\Delta_{\mathscr{A}}\Delta_{\mathscr{R}}+\Delta_{\mathscr{R}}^{2})$ . Then,
$\sigma(T)\leqslant 2n^{2}-2(\Delta_{\mathscr{A}}\Delta_{\mathscr{R}})n+2(\Delta_{\mathscr{A}}^{2}-\Delta_{\mathscr{A}}\Delta_{\mathscr{R}}+\Delta_{\mathscr{R}}^{2}).$ (10)
Therefore, the relationship between Sigma index and the Albertson index holds $\sigma(T)\geqslant\Delta_{\mathscr{A}}\operatorname{irr}(T)+2(\Delta_{\mathscr{A}}\Delta_{\mathscr{R}})n$. Then,
$\sigma(T)\leqslant 2n^{2}+\sum_{v\in V(T)}\deg_{T}(v)^{3}+\operatorname{irr}(T),$ (11)
from (9) and (11) yields
$\sigma(T)\leqslant\sum_{v\in V(T)}\deg_{T}(v)^{3}+\operatorname{irr}(T)+\left\lfloor\frac{n-2}{a_{r}-t_{m}}\right\rfloor.$ (12)
Finally, from (11) and (12) we find that the upper bound of Sigma index satisfy by (8). ∎
@@ 3.1 Study Bounds of Sigma Index Among Albertson Index
The following Lemma 3.1 states that, given specific constraints on the degree sequence $\mathscr{D}$ and order of a graph, the Sigma index exceeds the irregularity index by a strictly positive quantity consisting of explicit additive components with condition $n\leqslant\Delta_{\mathscr{A}}(a_{r}-a_{1})+\Delta_{\mathscr{R}}(t_{m}-t_{1})<\operatorname{irr}(T)$.
@@
Let $T$ be tree of order $n$, $\Delta_{\mathscr{A}}$ be the maximum degree of $\mathscr{A}$ and $\Delta_{\mathscr{R}}$ be the maximum degree of $\mathscr{R}$. If $n\leqslant\Delta_{\mathscr{A}}(a_{r}-a_{1})+\Delta_{\mathscr{R}}(t_{m}-t_{1})<\operatorname{irr}(T)$. Then, the lower bound of Sigma index satisfy
$\sigma(T)\geqslant\operatorname{irr}(T)+\dfrac{\left\lfloor\dfrac{2n}{a_{r}-a_{1}}\right\rfloor+\left\lceil\dfrac{2m}{n}\right\rceil}{n}+4n\Delta.$ (13)
@@
Recall $\mathscr{D}=(d_{1},d_{2},\dots,d_{n-1},d_{n})$ be a degree sequence where $d_{n}\geqslant d_{n-1}\geqslant\dots\geqslant d_{2}\geqslant d_{1}$, and $\mathscr{A}=(a_{1},a_{2},\dots,a_{r})$ where $a_{1}=(d_{2}+d_{1})/2,a_{2}=(d_{3}+d_{2})/2,\dots,a_{r}=(d_{n}+d_{n-1})/2$. Then, we noticed that $\Delta>a_{r}-a_{1}$ and $\sigma(T)>\operatorname{irr}(T)+\Delta_{\mathscr{A}}(a_{r}-a_{1})$. Then, we consider
$\operatorname{irr}(T)\leqslant\left\lfloor\frac{2n^{2}}{a_{r}-a_{1}}\right\rfloor.$ (14)
@@
$\sigma(T)-\operatorname{irr}(T)\geqslant\left\lfloor\frac{2n^{2}}{a_{r}-a_{1}}\right\rfloor+4\left\lceil\dfrac{2m}{n}\right\rceil\Delta.$ (15)
Thus, by considering $2m/n<n$ and $4n>\Delta$ we find that $\sigma(T)<4n^{2}\Delta$. Thus, according to (15) noticed that $\Delta>a_{r}-a_{1}$ and $n>a_{r}-a_{1}$. Then,
$\sigma(T)\geqslant\operatorname{irr}(T)+\left\lfloor\frac{2n^{2}}{a_{r}-a_{1}}\right\rfloor+\frac{4n^{2}\Delta}{a_{r}-a_{1}}.$ (16)
@@
$\sigma(T)\geqslant\dfrac{\left\lfloor\dfrac{2n}{a_{r}-a_{1}}\right\rfloor+\left\lceil\dfrac{2m}{n}\right\rceil}{n}+4n\Delta,\quad\operatorname{irr}(T)\geqslant\frac{n}{a_{r}-a_{1}}+n,$
@@
Lemma 3.1 frequently arises in the context of extremal graph theory and the analysis of topological indices (the Albertson and Sigma indices). This lemma is often employed to establish inequalities (as we show that among Figure 1) or bounds in Proposition 3.3 and 3.4 on these indices by utilizing degree sequences and the maximum vertex degree.
@@
Let $T$ be tree of order $n$, $\Delta_{\mathscr{A}}$ be the maximum degree of $\mathscr{A}$ and $\Delta_{\mathscr{R}}$ be the maximum degree of $\mathscr{R}$. Then, the lower bound of Sigma index satisfy
$\sigma(T)\geqslant\sqrt{\lambda_{\mathscr{D}}\sum_{i=1}^{n}d_{i}^{3}}-\left(\left\lfloor\frac{2n}{\lambda_{\mathscr{A}}}\right\rfloor+\left\lceil\frac{2m}{\lambda_{\mathscr{R}}}\right\rceil\right)+(n-\Delta)^{2}.$ (17)
@@
$\sigma(T)\leqslant 2n^{2}+\sum_{v\in V(T)}\deg_{T}(v)^{3}.$
Thus, when $\lambda_{\mathscr{D}}<n$ and $\lambda_{\mathscr{D}}\leqslant\Delta$. Also, for $\lambda_{\mathscr{A}}$ and $\lambda_{\mathscr{R}}$ satisfied. Thus, we need to compare the tem $2n/\lambda_{\mathscr{A}}$ and the term $2m/\lambda_{\mathscr{R}}$ with the term $2n/\lambda_{\mathscr{D}}$ by considering the term $2n/\lambda_{\mathscr{D}}$ is constant term (by examining values we find that the term equals 14). Thus,
@@
Therefore, from (18) by considering $(n-\Delta)^{2}$ is growing up and for Sigma index satisfy $\sigma(T)>(n-\Delta)^{2}$. Then, we find that $\sigma(T)>(n-\Delta)^{2}+\frac{2n}{\lambda_{\mathscr{D}}}$. Thus, we have
$\sigma(T)\geqslant(n-\Delta)^{2}+\frac{2n}{\lambda_{\mathscr{D}}}+\left\lfloor\frac{2n}{\lambda_{\mathscr{A}}}\right\rfloor+\left\lceil\frac{2m}{\lambda_{\mathscr{R}}}\right\rceil.$ (19)
@@
$\sigma(T)\geqslant(n-\Delta)^{2}+\sqrt{\lambda_{\mathscr{D}}\sum_{i=1}^{n}d_{i}^{3}}.$ (20)
@@
$\left(\left\lfloor\frac{2n}{\lambda_{\mathscr{A}}}\right\rfloor+\left\lceil\frac{2m}{\lambda_{\mathscr{R}}}\right\rceil\right)-\frac{2n}{\lambda_{\mathscr{D}}}\leqslant\sqrt{\lambda_{\mathscr{D}}\sigma(T)}.$
@@
Among Lemma 3.2, we provide the lower bound combines the terms of both $\lambda_{\mathscr{D}}$, $\lambda_{\mathscr{A}}$ and $\lambda_{\mathscr{R}}$, where $2n^{2}>\lambda_{\mathscr{D}}$ and $\sqrt{\lambda_{\mathscr{D}}n^{3}}\leqslant(n-\Delta)^{2}$. Lemma 3.2 emphasizes a stronger dependence on $\lambda_{\mathscr{D}}$ where $\left\lfloor 3\Delta t/2\right\rfloor\leqslant 3t(n-1)/2$ by considering the Albertson index $\operatorname{irr}(T)$ to measure degree heterogeneity.
@@
Let $T$ be tree of order $n$, $\Delta_{\mathscr{A}}$ be the maximum degree of $\mathscr{A}$ and $\Delta_{\mathscr{R}}$ be the maximum degree of $\mathscr{R}$. Then, the lower bound of Sigma index satisfy
$\sigma(T)\geqslant\frac{1}{3}\lambda_{\mathscr{D}}^{2}\left(\left\lfloor\dfrac{3n+1}{2}\right\rfloor+\left\lceil\dfrac{3m+1}{2}\right\rceil+\left\lfloor\dfrac{3\Delta+2n}{4}\right\rfloor\right)-\sum_{v\in V(T)}\deg_{T}(v)^{3}+\operatorname{irr}(T).$ (21)
@@
$t\left\lfloor\dfrac{3n+1}{2}\right\rfloor+2t\left\lceil\dfrac{3n-2}{2}\right\rceil+\left\lfloor\dfrac{3\Delta t}{2}\right\rfloor\leqslant\sigma(T).$ (23)
@@
$t\left\lfloor\dfrac{6k+1}{2}\right\rfloor+2t\left\lceil 3k-1\right\rceil+\left\lfloor\dfrac{3\Delta t}{2}\right\rfloor\leqslant\sigma(T).$ (24)
@@
Therefore, according to (26) for the term $t/2\leqslant k$. Then, $\sigma(T)\geqslant 12tk-t/2$. Thus, the relationship (24) holds.
@@
$t\left\lfloor 3k+2\right\rfloor+2t\left\lceil\dfrac{6k+1}{2}\right\rceil+\left\lfloor\dfrac{3\Delta t}{2}\right\rfloor\leqslant\sigma(T).$ (27)
@@
$\sigma(T)\leqslant\frac{1}{3}\lambda_{\mathscr{D}}^{2}\left(\left\lfloor\dfrac{3n+1}{2}\right\rfloor+\left\lceil\dfrac{3m+1}{2}\right\rceil+\left\lfloor\dfrac{3\Delta+2n}{4}\right\rfloor\right).$ (30)
Therefore, Sigma index satisfy $\sigma(T)\leqslant\operatorname{irr}(T)+n^{3}$ and when $n^{3}\leqslant\sum_{v\in V(T)}\deg_{T}(v)^{3}$ we find that (21) holds and satisfy the lower bound of Sigma index. ∎ Furthermore, we presented the following Table 1, to emphasize that results among Lemma 3.3, where $T_{1}$ and $T_{2}$ given as
@@
The positions labeled $n,\sigma,\operatorname{irr}(T),T_{1}$ and $T_{2}$ correspond to the variables whose pairwise correlation coefficients are displayed in the correlation matrix as
@@
Noticed that, regression coefficients [94764.10761811, 27387.84112146][\,-94764.10761811\,,\,27387.84112146\,], intercept 506577.67445476603-506577.67445476603, Model $R^{2}$ is $0.864810869832136$ and predicted Sigma for $n=350$ is 32304623.284719232-32304623.284719232.
Through the following Figure 2, we present analyze the data among Table 1 to consider the results in Lemma 3.3 cleared.
@@
4 Bounds of Sigma Index

### v3

@@ 3.1 Bounds on the Sigma Index Using Auxiliary Sequences
Let $\mathcal{CT}$ be a caterpillar tree with degree sequence $\mathscr{D}$. Then,
$\sigma(\mathcal{CT})\geqslant\operatorname{irr}(\mathcal{CT})+\left\lfloor\frac{n-2}{\Delta_{\mathscr{A}}-\Delta_{\mathscr{R}}}\right\rfloor+\Delta(\Delta_{\mathscr{A}}-\Delta_{\mathscr{R}})^{2},$ (5)
where $\Delta=\max\mathscr{D}$ is the maximum degree.
@@
[…]
$(d_{w_{i}}-1)^{2}-(d_{w_{i}}-1)=(d_{w_{i}}-1)(d_{w_{i}}-2).$
Thus, we noticed that over all pendants (let $k$ be the number of pendants, $k\approx n$- spine length), and bounding backbone contributions minimally ($\geq 0$), we obtain a lower estimate. The term $\left\lfloor(n-2)/(\Delta_{\mathscr{A}}-\Delta_{\mathscr{R}})\right\rfloor$ arises from the pendants supportable before the average/spread limits irregularity growth (since $\Delta_{\mathscr{A}}-\Delta_{\mathscr{R}}$ bounds consecutive variation). The quadratic correction $\Delta(\Delta_{\mathscr{A}}-\Delta_{\mathscr{R}})^{2}$ accounts for worst-case amplification when high-degree vertices concentrate differences. Combining yields the stated bound. ∎
@@
Under the same assumptions,
$\sigma(\mathcal{CT})\leqslant\sum_{v\in V(\mathcal{CT})}d_{\mathcal{CT}}(v)^{3}+\operatorname{irr}(\mathcal{CT})+\left\lfloor\frac{n-2}{\Delta_{\mathscr{A}}-\Delta_{\mathscr{R}}}\right\rfloor+\Delta(\Delta_{\mathscr{A}}-\Delta_{\mathscr{R}})^{2}.$ (6)
@@
Consider $\sigma(G)\leq\sum_{v\in V(\mathcal{CT})}d_{\mathcal{CT}}(v)^{3}-4m^{2}+O(\operatorname{irr}(G))$, with the cubic term dominating in high-degree cases. Adding the lower-bound correction from Proposition 3.3 and bounding residual terms via the sequence spread $(\Delta_{\mathscr{A}}-\Delta_{\mathscr{R}})$ yields the upper estimate. ∎
The proofs of Propositions 3.3–3.5 (assumed to follow similar patterns for refined cases, e.g., uniform pendants or specific $\Delta$) extend these ideas.
Collectively, it establishes that the Sigma index $\sigma(\mathcal{CT})$ is bounded above and below in terms of the Albertson index $\operatorname{irr}(\mathcal{CT})$ plus correction terms that depend explicitly on the auxiliary maximum values $\Delta_{\mathscr{A}}$ and $\Delta_{\mathscr{R}}$, as well as the effective transition count bounded by $\lfloor(n-2)/(\Delta_{\mathscr{A}}-\Delta_{\mathscr{R}})\rfloor$. This shows that $\sigma(\mathcal{CT})$ amplifies irregularity quadratically through degree differences (via the $(\Delta_{\mathscr{A}}-\Delta_{\mathscr{R}})^{2}$ term and cubic degree contributions), while $\operatorname{irr}(\mathcal{CT})$ grows only linearly. In caterpillar trees—often extremal for irregularity under fixed degree sequences—these bounds highlight how sequence smoothness (small $\Delta_{\mathscr{A}}-\Delta_{\mathscr{R}}$) forces tighter irregularity, whereas large spreads allow significantly higher $\sigma$ values.
@@ 3.2 Study Bounds of Sigma Index Among Albertson Index
The following Lemma 3.1 states that, given specific constraints on the degree sequence $\mathscr{D}$ and order of a graph, the Sigma index exceeds the irregularity index by a strictly positive quantity consisting of explicit additive components with condition $n\leqslant\Delta_{\mathscr{A}}(a_{r}-a_{1})+\Delta_{\mathscr{R}}(t_{m}-t_{1})<\operatorname{irr}(\mathcal{CT})$.
@@
Consider $\mathcal{CT}$ be a caterpillar tree, $\Delta_{\mathscr{A}}$ be the maximum degree of $\mathscr{A}$ and $\Delta_{\mathscr{R}}$ be the maximum degree of $\mathscr{R}$. If $n\leqslant\Delta_{\mathscr{A}}(a_{r}-a_{1})+\Delta_{\mathscr{R}}(t_{m}-t_{1})<\operatorname{irr}(\mathcal{CT})$. Then, the lower bound of Sigma index satisfy
$\sigma(\mathcal{CT})\geqslant\operatorname{irr}(T)+\dfrac{\left\lfloor\dfrac{2n}{a_{r}-a_{1}}\right\rfloor+\left\lceil\dfrac{2m}{n}\right\rceil}{n}+4n\Delta.$ (7)
@@
Recall $\mathscr{D}=(d_{1},d_{2},\dots,d_{n-1},d_{n})$ be a degree sequence where $d_{n}\geqslant d_{n-1}\geqslant\dots\geqslant d_{2}\geqslant d_{1}$, and $\mathscr{A}=(a_{1},a_{2},\dots,a_{r})$ where $a_{1}=(d_{2}+d_{1})/2,a_{2}=(d_{3}+d_{2})/2,\dots,a_{r}=(d_{n}+d_{n-1})/2$. Then, we noticed that $\Delta>a_{r}-a_{1}$ and $\sigma(\mathcal{CT})>\operatorname{irr}(\mathcal{CT})+\Delta_{\mathscr{A}}(a_{r}-a_{1})$. Then, we consider
$\operatorname{irr}(\mathcal{CT})\leqslant\left\lfloor\frac{2n^{2}}{a_{r}-a_{1}}\right\rfloor.$ (8)
@@
$\sigma(\mathcal{CT})-\operatorname{irr}(\mathcal{CT})\geqslant\left\lfloor\frac{2n^{2}}{a_{r}-a_{1}}\right\rfloor+4\left\lceil\dfrac{2m}{n}\right\rceil\Delta.$ (9)
Thus, by considering $2m/n<n$ and $4n>\Delta$ we find that $\sigma(\mathcal{CT})<4n^{2}\Delta$. Thus, according to (9) noticed that $\Delta>a_{r}-a_{1}$ and $n>a_{r}-a_{1}$. Then,
$\sigma(\mathcal{CT})\geqslant\operatorname{irr}(\mathcal{CT})+\left\lfloor\frac{2n^{2}}{a_{r}-a_{1}}\right\rfloor+\frac{4n^{2}\Delta}{a_{r}-a_{1}}.$ (10)
@@
$\sigma(\mathcal{CT})\geqslant\dfrac{\left\lfloor\dfrac{2n}{a_{r}-a_{1}}\right\rfloor+\left\lceil\dfrac{2m}{n}\right\rceil}{n}+4n\Delta,\quad\operatorname{irr}(\mathcal{CT})\geqslant\frac{n}{a_{r}-a_{1}}+n,$
@@
Lemma 3.1 frequently arises in the context of extremal graph theory and the analysis of topological indices (the Albertson and Sigma indices). This lemma is often employed to establish inequalities or bounds in Proposition 3.3 and 3.4 on these indices by utilizing degree sequences and the maximum vertex degree.
@@
Consider $\mathcal{CT}$ be a caterpillar tree, $\Delta_{\mathscr{A}}$ be the maximum degree of $\mathscr{A}$ and $\Delta_{\mathscr{R}}$ be the maximum degree of $\mathscr{R}$. Then, the lower bound of Sigma index satisfy
$\sigma(\mathcal{CT})\geqslant\sqrt{\lambda_{\mathscr{D}}\sum_{i=1}^{n}d_{i}^{3}}-\left(\left\lfloor\frac{2n}{\lambda_{\mathscr{A}}}\right\rfloor+\left\lceil\frac{2m}{\lambda_{\mathscr{R}}}\right\rceil\right)+(n-\Delta)^{2}.$ (11)
@@
$\sigma(\mathcal{CT})\leqslant 2n^{2}+\sum_{v\in V(\mathcal{CT})}d_{C}C(v)^{3}.$
Thus, when $\lambda_{\mathscr{D}}<n$ and $\lambda_{\mathscr{D}}\leqslant\Delta$. Also, for $\lambda_{\mathscr{A}}$ and $\lambda_{\mathscr{R}}$ satisfied. Thus, we need to compare the term $2n/\lambda_{\mathscr{A}}$ and the term $2m/\lambda_{\mathscr{R}}$ with the term $2n/\lambda_{\mathscr{D}}$ by considering the term $2n/\lambda_{\mathscr{D}}$ is constant term (by examining values we find that the term equals 14). Thus,
@@
Therefore, from (12) by considering $(n-\Delta)^{2}$ is growing up and for Sigma index satisfy $\sigma(\mathcal{CT})>(n-\Delta)^{2}$. Then, we find that $\sigma(\mathcal{CT})>(n-\Delta)^{2}+\frac{2n}{\lambda_{\mathscr{D}}}$. Thus, we have
$\sigma(\mathcal{CT})\geqslant(n-\Delta)^{2}+\frac{2n}{\lambda_{\mathscr{D}}}+\left\lfloor\frac{2n}{\lambda_{\mathscr{A}}}\right\rfloor+\left\lceil\frac{2m}{\lambda_{\mathscr{R}}}\right\rceil.$ (13)
@@
$\sigma(\mathcal{CT})\geqslant(n-\Delta)^{2}+\sqrt{\lambda_{\mathscr{D}}\sum_{i=1}^{n}d_{i}^{3}}.$ (14)
@@
$\left(\left\lfloor\frac{2n}{\lambda_{\mathscr{A}}}\right\rfloor+\left\lceil\frac{2m}{\lambda_{\mathscr{R}}}\right\rceil\right)-\frac{2n}{\lambda_{\mathscr{D}}}\leqslant\sqrt{\lambda_{\mathscr{D}}\sigma(\mathcal{CT})}.$
@@
Among Lemma 3.2, we provide the lower bound combines the terms of both $\lambda_{\mathscr{D}}$, $\lambda_{\mathscr{A}}$ and $\lambda_{\mathscr{R}}$, where $2n^{2}>\lambda_{\mathscr{D}}$ and $\sqrt{\lambda_{\mathscr{D}}n^{3}}\leqslant(n-\Delta)^{2}$. Lemma 3.2 emphasizes a stronger dependence on $\lambda_{\mathscr{D}}$ where $\left\lfloor 3\Delta t/2\right\rfloor\leqslant 3t(n-1)/2$ by considering the Albertson index $\operatorname{irr}(\mathcal{CT})$ to measure degree heterogeneity.
@@
Let $\mathcal{CT}$ be a caterpillar tree, $\Delta_{\mathscr{A}}$ be the maximum degree of $\mathscr{A}$ and $\Delta_{\mathscr{R}}$ be the maximum degree of $\mathscr{R}$. Then, the lower bound of Sigma index satisfy
$\sigma(\mathcal{CT})\geqslant\frac{1}{3}\lambda_{\mathscr{D}}^{2}\left(\left\lfloor\dfrac{3n+1}{2}\right\rfloor+\left\lceil\dfrac{3m+1}{2}\right\rceil+\left\lfloor\dfrac{3\Delta+2n}{4}\right\rfloor\right)-\sum_{v\in V(\mathcal{CT})}d_{\mathcal{CT}}(v)^{3}+\operatorname{irr}(\mathcal{CT}).$ (15)
@@
$t\left\lfloor\dfrac{3n+1}{2}\right\rfloor+2t\left\lceil\dfrac{3n-2}{2}\right\rceil+\left\lfloor\dfrac{3\Delta t}{2}\right\rfloor\leqslant\sigma(\mathcal{CT}).$ (17)
@@
$t\left\lfloor\dfrac{6k+1}{2}\right\rfloor+2t\left\lceil 3k-1\right\rceil+\left\lfloor\dfrac{3\Delta t}{2}\right\rfloor\leqslant\sigma(\mathcal{CT}).$ (18)
@@
Therefore, according to (20) for the term $t/2\leqslant k$. Then, $\sigma(\mathcal{CT})\geqslant 12tk-t/2$. Thus, the relationship (18) holds.
@@
$t\left\lfloor 3k+2\right\rfloor+2t\left\lceil\dfrac{6k+1}{2}\right\rceil+\left\lfloor\dfrac{3\Delta t}{2}\right\rfloor\leqslant\sigma(\mathcal{CT}).$ (21)
@@
$\sigma(\mathcal{CT})\leqslant\frac{1}{3}\lambda_{\mathscr{D}}^{2}\left(\left\lfloor\dfrac{3n+1}{2}\right\rfloor+\left\lceil\dfrac{3m+1}{2}\right\rceil+\left\lfloor\dfrac{3\Delta+2n}{4}\right\rfloor\right).$ (24)
Therefore, Sigma index satisfy $\sigma(\mathcal{CT})\leqslant\operatorname{irr}(\mathcal{CT})+n^{3}$ and when $n^{3}\leqslant\sum_{v\in V(\mathcal{CT})}d_{\mathcal{CT}}(v)^{3}$ we find that (15) holds and satisfy the lower bound of Sigma index. ∎ Furthermore, we presented the following Table 1, to emphasize that results among Lemma 3.3, where $T_{1}$ and $T_{2}$ given as
@@
The positions labeled $n,\sigma,\operatorname{irr}(\mathcal{CT}),T_{1}$ and $T_{2}$ correspond to the variables whose pairwise correlation coefficients are displayed in the correlation matrix as
@@
The correlation matrix $M_{\mathscr{D}}$ reveals extremely strong positive linear associations among the considered variables. In particular:
  • $\sigma(T)$ exhibits a very high correlation with $T_{2}$ (coefficient $0.993$), indicating that the composite term $T_{2}=\lambda_{\mathscr{D}}^{2}T_{1}$ serves as an excellent predictor of the Sigma index.
  • $\sigma(T)$ also correlates strongly with $\operatorname{irr}(T)$ (coefficient $0.930$), consistent with the quadratic amplification of degree differences in the Sigma index compared to the linear nature of the Albertson index.
  • Near-perfect correlations (close to $1$) appear between size-related parameters (e.g., $T_{1}$) and other quantities, reflecting the scaling behavior of irregularity measures with graph order and degree constraints.
A linear regression of $\sigma(T)$ against $T_{2}$ yields a slope of approximately $0.970$, an intercept near 857-857, and an $R^{2}$ value of essentially $1.000$ (to numerical precision). This near-perfect fit demonstrates that $\sigma(T)\approx 0.970\cdot T_{2}-857$ holds very closely for the tested trees, confirming that the lower bound expression derived in Lemma 3.3 is tight or nearly attained in these instances.
Figure 2 visualizes the relationship between $\sigma(T)$ and $T_{2}$. The points lie almost exactly on the fitted regression line, providing strong empirical support for the analytical lower bound and highlighting the effectiveness of the proposed proxy $T_{2}$ (which embeds the quadratic dependence on the average degree $\lambda_{\mathscr{D}}$) in capturing the behavior of the Sigma index.
Figure 2: Sigma index $\sigma(T)$ versus the lower-bound proxy $T_{2}$ for the trees in Table 1. The near-perfect linear fit confirms that $\sigma(T)$ closely attains the proposed lower bound.
These computational results reinforce the theoretical findings: the Sigma index grows quadratically with parameters derived from the degree sequence (via $\lambda_{\mathscr{D}}^{2}$), while remaining closely aligned with the suggested lower-bound form. This close correspondence validates the extremal analysis and suggests that caterpillar (or near-caterpillar) realizations with the given degree sequences achieve values very close to the theoretical minimum for $\sigma(T)$. The high correlations and excellent fit further motivate focusing subsequent bounds and extremal constructions on similar composite terms involving average degree squared, maximum degree, and graph order. Through the following Figure 3, we present analyze the data among Table 1 to consider the results in Lemma 3.3 cleared.
@@
4 Extremal Bounds on the Sigma Index in Caterpillar Trees